Please use financial calculator to assist in showing the work.

A project's initial cost is $8,500. The project will produce cash inflows of $4,200 a year for 4 years with a final cash inflow of $6,700 in year 5. What is the net present value of this project if the required rate of return is 15 percent?
Answer #1

answer:- NET PRESENT VALUE =$6821.99
Answer #2

Step-by-Step Calculation (Using Financial Calculator Logic):

  1. Initial Cost (Outflow):

    • Today (Year 0): $8,500 (negative because it’s an expense).

  2. Annual Cash Inflows (Years 1–4):

    • PMT = $4,200, N = 4, I/Y = 15%.

    • $4,200 each year for 4 years.


    • Calculator Input:

    • Present Value (PV) of Annuity:

      PVYears 1–4=4,200×(1(1+0.15)40.15)=4,200×2.8550=$11,991.


  3. Final Cash Inflow (Year 5):

    • $6,700 (one-time payment in Year 5).

    • Present Value (PV) of Single Sum:

      PVYear 5=6,700(1+0.15)5=6,7002.0114=$3,331.

  4. Total Present Value of Inflows:

    PVTotal Inflows=PVYears 1–4+PVYear 5=$11,991+$3,331=$15,322.

  5. Net Present Value (NPV):

    NPV=PVTotal InflowsInitial Cost=$15,322$8,500=$6,822.


Answer:
             The correct NPV is $6,822.


answered by: anonymous